# Sommerfeld Enhancement for Vector Dark Matter in the Fundamental Representation of $SU(2)_{L}$

### v1

@@ Acknowledgments
This work was partially found by Fondecyt grant 1230110, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044. We are very thankful to Motoko Fujiwara for discussions and suggestions that contributed significantly to improve the quality of this paper.
@@ A.1 Gauge-Dark Matter
$J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})=2e^{2}s_{W}^{2}(c_{W}-s_{W}t_{W}^{2})$ $(-[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})]$ (118)
$J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})$ $=-g^{2}s_{W}^{2}\theta_{W}[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})$
@@
$J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})=-\dfrac{g}{2}\left[v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})-v_{-}^{\dagger}(x)v_{+}^{\dagger}(x^{0},\vec{y})a_{0}(x)v_{0}(x^{0},\vec{y})\right]$ (120)
@@
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{Gauge}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (121)
where $V_{Gauge}(|\vec{x}-\vec{y}|)$ is the electroweak matrix, with $r=|\vec{x}-\vec{y}|$, therefore:
$V_{Gauge}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4s_{W}^{2}e^{-M_{Z}r}&0&0&0\\ 0&4s_{W}^{2}e^{-M_{Z}r}&0&0\\ 0&0&16s_{W}^{2}e^{-M_{Z}r}&2\sqrt{2}e^{-M_{W}r}\\ 0&0&2\sqrt{2}e^{-M_{W}r}&8s_{W}^{2}e^{-M_{Z}r}-2(c_{W}-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (126)
@@ A.2 Higgs Boson-Dark Matter
$V_{vector}(r)=\dfrac{v^{2}}{32\pi m^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (132)
this is electroweak matrix (Higgs Boson-Dark Matter interaction)
@@ B.1 Gauge-Dark Matter Interaction
The non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2){L}$. In the framework of effective quantum field theory, these annihilations are depicted by the inclusion of an imaginary potential or an absorptive term in the action, which accounts for the disappearance of pairs of non-relativistic particles. These absorptive terms originate from $S{eff,2}$ as outlined in equation (56). To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams. In this representation, the matrix $\Delta_{0}^{-1}$ characterizes the propagators of gauge fields, and each $\mathcal{Q}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al. [7].
@@
Therefore, we perform an integration over the timelike component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell. The integration of 4-momentum is then expressed as:
@@
with add angular momentun matrices
@@
The annilihation matrices (Gauge-Dark Matter) are:
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi m^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (180)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (185)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (190)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (195)
@@
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (200)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (205)
@@
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi m^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (210)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi m^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (215)
@@ B.2 Higgs-Dark Matter Interaction
$\Gamma_{vector}=\dfrac{1}{32\pi m^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (221)

### v2

@@ Acknowledgments
This work was partially found by Fondecyt grant 1230110, ANID PIA/APOYO AFB220004, and ANID—The Millennium Science Initiative Program ICN2019_044. We are very thankful to Motoko Fujiwara and Patricio Escalona for discussions and suggestions that contributed significantly to improve the quality of this paper.
@@ A.1 Gauge-Dark Matter
$J_{A}^{0}(x)J_{A}^{0}(x^{0},\vec{y})=-2g^{2}(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}$ $(-[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})]$ (56)
$J_{Z}^{0}(x)J_{Z}^{0}(x^{0},\vec{y})$ $=-g^{2}(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})^{2}[v_{+}^{\dagger}(x)v_{-}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})$
@@
$J_{W_{+}}^{0}(x)J_{W_{-}}^{0}(x^{0},\vec{y})=$ $-\left[\left(\dfrac{3}{8}-2\sqrt{2}\right)(\kappa_{1}+\kappa_{2})+2\sqrt{2}\right]^{2}$
$g^{2}\left[v_{0}^{\dagger}(x)a_{0}^{\dagger}(x^{0},\vec{y})v_{+}(x)v_{-}(x^{0},\vec{y})-v_{-}^{\dagger}(x)v_{+}^{\dagger}(x^{0},\vec{y})a_{0}(x)v_{0}(x^{0},\vec{y})\right]$ (58)
@@
$S_{eff,1_{NR}}=-\int d^{4}xd^{3}y~{}s(x,\vec{y})^{\dagger}V_{g-V}(|\vec{x}-\vec{y}|)s(x,\vec{y})$ (59)
where $V_{g-V}(|\vec{x}-\vec{y}|)$ is the electroweak matrix, with $r=|\vec{x}-\vec{y}|$, therefore:
$V_{g-V}(r)=-\dfrac{g^{2}}{8\pi r}\begin{pmatrix}4a^{2}e^{-M_{Z}r}&0&0&0\\ 0&4a^{2}e^{-M_{Z}r}&0&0\\ 0&0&16a^{2}e^{-M_{Z}r}&b^{2}e^{-M_{W}r}\\ 0&0&b^{2}e^{-M_{W}r}&a^{2}e^{-M_{Z}r}-2(c_{W}(1-(\kappa_{1}+\kappa{2})-s_{W}t_{W}^{2})^{2}\end{pmatrix}$ (60)
with $a=(c_{W}(\kappa_{1}+\kappa_{2})+s_{W})$, $b=(\frac{3}{8}-2\sqrt{2})(\kappa_{1}+\kappa_{2})+2\sqrt{2}$
@@ A.2 Higgs Boson-Dark Matter
$V_{h-V}(r)=\dfrac{v^{2}}{32\pi M_{V}^{2}r}\begin{pmatrix}8\lambda_{4}^{2}a&0&0&0\\ 0&8\lambda_{4}^{2}a&0&0\\ 0&0&\dfrac{1}{2}(\lambda_{2}+\lambda_{3})^{2}a&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}\\ 0&0&(\lambda_{3}+2\lambda_{4})^{2}e^{-M_{W}r}&\lambda_{2}^{2}a\end{pmatrix}$ (62)
with $a=(2e^{-M_{h}r}+e^{-M_{Z}r})$, this is electroweak matrix (Higgs Boson-Dark Matter interaction)
@@ B.1 Gauge-Dark Matter Interaction
The non-relativistic vector dark matter particles have the potential to annihilate into relativistic particles, specifically gauge bosons of $SU(2){L}$. In the framework of effective quantum field theory, these annihilations are depicted by the inclusion of an imaginary potential or an absorptive term in the action, which accounts for the disappearance of pairs of non-relativistic particles. These absorptive terms originate from $S_{eff,2}$ as outlined in equation (38). To provide an interpretation, we can visualize $S_{eff,2}$ in terms of Feynman diagrams. In this representation, the matrix $\Delta_{0}^{-1}$ characterizes the propagators of gauge fields, and each $\mathcal{Q}$ represents a vertex featuring two gauge fields and two vector dark matter fields, as described by Chowdhury et al. [4].
@@
Therefore, we perform an integration over the time-like component of the 4-momentum, resulting in the transformation $d^{4}q\delta(q^{2})\rightarrow\frac{d^{3}q}{2|\vec{q}|}$ when the gauge particles are on-shell. The integration of 4-momentum is then expressed as:
@@
with add angular momentum matrices
@@
The annihilation matrices (Gauge-Dark Matter) are:
$\Gamma_{WW}^{J=0}=\dfrac{2g^{4}}{64\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (78)
$\Gamma_{WW}^{J=2}=\dfrac{g^{4}}{64\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2&0&0&0\\ 0&2&0&0\\ 0&0&1&1\\ 0&0&1&1\end{pmatrix}$ (79)
$\Gamma_{AA}^{J=0}=\dfrac{e^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (80)
$\Gamma_{AA}^{J=2}=\dfrac{e^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2a^{2}&0&0&0\\ 0&2a^{2}&0&0\\ 0&0&a^{2}&4a^{2}\\ 0&0&4a^{2}&4\end{pmatrix}$ (81)
@@
$\Gamma_{ZZ}^{J=0}$ $=\dfrac{g^{4}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$ (82)
$\Gamma_{ZZ}^{J=2}$ $=\dfrac{g^{4}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}2b^{2}&0&0&0\\ 0&2b^{2}&0&0\\ 0&0&2b^{2}&-4c\\ 0&0&-4c&4\sin^{4}\theta_{W}\end{pmatrix}$
@@
$\Gamma_{AZ}^{J=0}$ $=\dfrac{e^{2}g^{2}}{128\sqrt{3}\pi M_{V}^{2}}\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$ (83)
$\Gamma_{AZ}^{J=2}$ $=\dfrac{e^{2}g^{2}}{256\pi M_{V}^{2}}\left(\dfrac{1+3\cos(2\theta)}{2\sqrt{6}}+\sin^{2}\theta\right)\begin{pmatrix}d^{2}&0&0&0\\ 0&d^{2}&0&0\\ 0&0&d^{2}&f\\ 0&0&f&\dfrac{\sin^{2}(2\theta_{W})}{\cos^{2}\theta_{W}}\end{pmatrix}$
@@ B.2 Higgs-Dark Matter Interaction
$\Gamma_{h-V}=\dfrac{1}{32\pi M_{V}^{2}}\begin{pmatrix}\left(\dfrac{3}{2}+\sqrt{2}\right)^{2}\lambda_{4}^{2}&0&0&0\\ 0&\dfrac{9}{4}\lambda_{4}^{2}&0&0\\ 0&0&\dfrac{1}{4}(7\lambda_{2}^{2}+6\lambda_{2}\lambda_{3}+5\lambda_{3}^{2})&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}\\ 0&0&\dfrac{1}{2}(\lambda_{3}+2\lambda_{4})^{2}&\dfrac{3}{4}\lambda_{2}^{2}+(\lambda_{2}+\lambda_{3})^{2}\end{pmatrix}$ (85)